OpenStudy (nbarrera):

What is -5n+15=-125?

OpenStudy (anonymous):

-22

OpenStudy (anonymous):

28

OpenStudy (anonymous):

-5n +15 = -125 -15 = -15 -5n = 110 / -5 = /-5 n = -22

OpenStudy (coconutjj):